### v4

@@ 4 Data
We consider all 1,072 SPACs traded in major exchanges between January 2010 and December 2021. See Figure A4 for the monthly scatter plot in Figure A4, which includes the so-called boom and bust cycle for SPACs. We construct our sample by merging five datasets: (i) the SPAC Research,555https://www.spacresearch.com/. (ii) the PrivateRaise–SPACs Platform,666https://www.privateraise.com/about/platform_spac.php. (iii) the PrivateRaise–PIPEs Platform (iv) Bloomberg data on SPAC prices, and (v) Super 8-K filings from SEC.777https://www.sec.gov/edgar/search-and-access.888Thus, we exclude SPACs traded in the over-the-counter markets. After the merger between a SPAC and a target is complete, the SPAC is delisted. The new (merged) company may be assigned a new ticker symbol. We use Bloomberg-assigned unique identifiers to match across samples.
The SPAC Research database and PrivateRaise’s SPACs platform database contain information on SPACs, including important event dates. PrivateRaise’s PIPE platform database includes comprehensive information on PIPE transactions. The platform covers more than 20,000 placement profiles of PIPE transactions. It also provides contract terms, investment amounts, types of transactions, prices, identities, and allocations of the PIPE investors.
For these 1,072 SPACs, we manually merged the first two databases. The issuer’s name recorded in the PIPE transactions can be either that of SPAC or the post-merger company. In those cases, we use both names to match the SPAC sample with the PrivateRaise database. We also use the total PIPE amount from PrivateRaise to cross-check the match to ensure match accuracy. If there are inconsistencies, we use the SPAC’s SEC filings documents to record the correct amount. We then merge our sample with Bloomberg’s pricing data; warrant prices are unavailable in CRSP. Our matching process yields detailed information on transactions for all three types of PIPE investments for each SPAC in our sample.
@@
Table 2 provides summary data. On average, SPACs raise $272.42 million proceeds during their IPO, with a median of $240 million and 90th percentile at $440 million. A typical SPAC unit contains 0.44 warrants and 0.01 rights. During the IPO process, SPACs’ underwriters choose, on average, to exercise the 11.59% over-allotment option to meet the additional demand.
@@
Money Left on the Table. We use money left on the table (“MLOT”) to measure the profit accrued by PIPE investors. The MLOT denotes the difference between the closing-day price on the first trading day of the post-merger company and the offer price presented to PIPE investors by the SPAC, and this difference is multiplied by the number of securities held by the PIPE investors. This calculation considers both common shares and warrants acquired by the PIPE investors. To determine the rate of underpricing, we normalize the MLOT by PIPE investors’ total investment amount.999Although the contracts with PIPE investors are signed before the merger is closed, they represent a pre-commitment price, implying that the PIPE investors deploy their capital only at the merger’s closure. This arrangement mirrors a ”one-day” return that traditional IPO investors experience. We carry out the MLOT and underpricing calculations for all three forms of PIPE investors (pre, at, and post).
@@
Premium PIPE Investors. In total, there are 2,005 PIPE investors in our sample. Out of them, 76 invested in PIPE (pre) transaction, 249 in PIPE (post), and 1,826 invested in at least one PIPE (at) transaction, but only a few large institutional investors participated frequently and invested significantly more than others. To capture this distinction, we use k-means clustering on the total investments in all SPACs and the number of SPACs to classify PIPE investors into two groups. We refer to a subset of PIPE investors who frequently participate as “premium” PIPE investors and the rest as non-premium PIPE investors. There are 34 premium PIPE investors in our sample (Table 3).101010As a way to corroborate this classification, we counted the number of articles written about each of these 34 premium investors in The New York Times and found that these premium investors have the highest counts. For example, excluding research universities, e.g., Columbia University, and non-financial companies, e.g., General Motors and Microsoft Inc., top ten investors include Morgan Stanley, Fidelity, USB, Wells Fargo, Blackstone, Citadel, Bain Capital, BlackRock, T. Rowe Price, and SoftBank.
@@
This comparison is consistent with the quid pro quo arrangement where non-premium investors attain high returns on successful deals, offset by low returns on less successful ones. In contrast, premium investors see more steady returns due to their superior information. The difference between the MLOTs of premium and non-premium investors is further analyzed in Section 6.
@@ 5.1.1 Underpricing during SPAC IPO
We first examine underpricing during the SPAC IPO. Table 2 reports that the average equal-weighted first-day return of SPACs is 1.3%. Although our data start from 2010, most SPAC IPOs occurred between 2020 and 2021. In 2020, 248 SPACs went public, with an average equal-weighted first-day return of 1.3%; in 2021, 613 SPACs went public, and their average equal-weighted first-day return was 1.51%. Although these measures are statistically different from zero (p-values in both cases), they are small.111111The average equal-weighted first-day return for traditional IPOs was 41.6% in 2020 and 32% in 2021; see https://site.warrington.ufl.edu/ritter/files/IPOs-Underpricing.pdf (accessed 09/22/2022). In addition, SPACs almost always trade at $10 per share until the target announcement date. This uniformity is inconsistent with any information about sponsor quality revealed during the SPAC IPO.
@@ 5.1.2 Underpricing during de-SPAC
From Table 2, we see that, on average, a private company that uses the SPAC route to access the public market leaves $126.93 million MLOT to the sponsors and $201.22 million to the PIPE investors. Furthermore, we also see that PIPE investors who participate in different stages of SPAC are compensated differently. During the SPAC IPO, PIPE (pre) investors earn an average of $29.37 million through forward-purchase agreements, but PIPE (at) investors get, on average, a total of $193.07 million per deal. After the target is disclosed, PIPE (post) investors who participate receive an additional $8.07 million MLOT. Even among PIPE (at) investors, approximately one-third of the MLOT accrues to 34 premium investors. This difference suggests that the two types of investors have different business models, which we return to later in Section 6.
Overall, the SPAC shares offered to PIPE (at) investors, especially the premium PIPE investors, during the de-SPAC phase are underpriced. These results are consistent with Hypothesis 1 that PIPE (at) investors participating in the de-SPAC stage are privately informed about the quality of the target company.
5.2 Information Production
@@ 5.2.1 Redemption Rates
To test the first part of Hypothesis 1-1, we use the following regression:
where denotes the investment allocated to premium PIPE (at) investors as a fraction of the total IPO proceeds in SPAC , similarly for other allocations, and is a vector of covariates including total IPO proceeds and fixed effects that may vary across specifications. If premium investors generate value-relevant information, the higher the share allocated to premium investors, the lower the redemption rate, i.e., . For inference, we cluster the robust standard errors at the sponsor level to capture correlations within a sponsor but across investors and SPACs.
Table 4 provides evidence consistent with the hypothesis. In column (1), we include only the allocations to four types of PIPE investors. In column (2), we control for the size of the SPAC IPO and the liquidation risk of the SPAC when the target is announced using Days Remaining until the expiration. In column (3), we control for the quarter-fixed effects. In column (4), we control for the target company’s sector fixed effects to alleviate the problem that investors may have varying investment preferences across industries.
@@
Note: This table shows the results from estimating (1) using SPACs that went through an IPO from January 2010 to December 2021. The independent variable is the investment amount as a percentage of the SPAC’s IPO proceeds for all PIPE (pre) investors, premium PIPE (at) investors, non-premium PIPE (at) investors, and all PIPE (post) investors. Premium PIPE (at) investors are listed in Table 3. All other variables are defined in Table B1. All unbounded continuous variables are winsorized at their and percentile values. Robust standard errors are clustered at the sponsor level and are reported in parentheses. In column (5), we use 1,000 Bootstrap samples to estimate the standard errors.. , and denote p-values less than 0.1, 0.05, and 0.01, respectively.
We find that the allocations to premium PIPE (at) investors negatively correlate with the redemption rates, and this result is robust across all specifications. In particular, the estimates suggest that one standard deviation (32.85 pps) increase in premium PIPE investors’ participation is associated with a 7.03 pps decrease in the redemption rate, which equals 12.4% of the sample mean.
@@
The estimates with the additional controls are in Table 4 column (5). Comparing the estimates in columns (4) and (5), we can see that our result is qualitatively similar, i.e., premium investors provide certification roles. The estimates suggest that one standard deviation (32.85 pps) increase in premium PIPE (at) investors’ participation is associated with a 7.19 pps decrease in the redemption rate, which equals 12.7% of the sample mean.
We also find no association between non-premium PIPE (at) or PIPE (pre) allocations and redemption rates, as in Hypothesis 1-1 and Hypothesis 2-1. These results are consistent with the hypothesis that premium PIPE (at) investors are better informed about the deal quality than either the non-premium PIPE (at) investors or PIPE (pre).
The negative association between premium PIPE (at) allocation and redemption rates is consistent with premium investors producing information. Suppose not, and instead, suppose that sponsors underprice and distribute them to the premium investors in exchange for a promise to help them in the future on an unrelated SPAC. In that case, because the investors’ timing should be immaterial, we expect the allocations to PIPE (post) investors to correlate negatively with the redemption rate. However, as we see from Table 4, the two are positively correlated, further strengthening our finding.121212Relatedly, almost all PIPE (post) investors are non-premium investors. In fact, among the top 20 PIPE (post) investors by investment size, only Luxor Capital Group is a premium investor. Thus, it is reasonable to assume that sponsors include premium investors at the time of the announcement as certification and use non-premium investors to backfill the trust fund.
@@ 5.2.2 Announcement-day Return
Note: This table shows the results from estimating (1), announcement-day return as the dependent variable, using SPACs that went through an IPO from January 2010 to December 2021. The key independent variable is the investment amount as a percentage of the SPAC’s IPO proceeds for all PIPE (pre) investors, premium PIPE (at) investors, non-premium PIPE (at) investors, and all PIPE (post) investors. Premium PIPE (at) investors are listed in Table 3. All other variables are defined in Table B1. All unbounded continuous variables are winsorized at their and percentile values. Robust standard errors are clustered at the sponsor level and are reported in parentheses. In column (5), we use 1,000 Bootstrap samples to estimate the standard errors. , and denote p-values less than 0.1, 0.05, and 0.01, respectively.
@@
Likewise, the hypothesis predicts a non-positive market reaction to the announcement of PIPE (post) investments. To assess this claim, we use an event study approach. In particular, we estimate the cumulative abnormal returns (CAR) (Campbell, Lo and MacKinlay, 1997) associated with announcements about PIPE investors while also announcing the target. The estimation period is from the SPAC’s unit separation date to the event window.
@@
Note: This table reports the SPAC share’s cumulative average abnormal return in percentage for three different event windows around announcements of 81 PIPE (post) transactions. Here, CAPM stands for capital asset pricing model, and FF3 and FF5 stand for Fama-French three- and five-factor models, respectively. The cross-sectional standard deviations of the cumulative abnormal returns are reported in parentheses. , and denote p-values less than 0.1, 0.05, and 0.01, respectively.
5.3 Quid Pro Quo
Next, we test Hypothesis 2 that there is a quid pro quo arrangement between sponsors and non-premium PIPE (at) investors. Recall that under favoritism, the current MLOT should be negatively correlated with the past MLOT. In particular, if the sponsor allocates securities to non-premium investors as a favor, and if the past MLOT is high, the sponsor will likely ask this investor to finance a weak deal today. We verify that the probability that a sponsor allocates securities to such a “relationship” investor increases if the investor has higher MLOT from the past deals and the SPAC deal is cold. Lastly, we show that quid pro quo lowers the liquidation risks.
@@ 5.3.1 Mean-reverting MLOT
where MLOT is the MLOT that PIPE (at) investor gets in the current SPAC deal initialized by the sponsor , and is the MLOT from the most recent deal in the past. As before, includes several covariates, including fixed effects (see Table 9).
@@
Under information production, the deviation of lagged MLOT from its mean is either uncorrelated or positively correlated with the current MLOT; hence in (3). Under quid pro quo arrangements, however, the MLOT should increase if the past deal was lower than the average and vice versa; hence, . The results are in column (1) of Table 9. As expected under favoritism, MLOT exhibits mean reversion.
@@
| (lag-MLOT-)(lag-MLOT-) | 121.96 | 85.54 | |
| (lag-MLOT-) | 135.78 | 5.77 | |
| (lag-MLOT-) | 166.93 | 297.54 | |
Note: This table shows results from estimating (4) using SPACs that went through IPO between January 2010 to December 2021. Each observation is at (SPAC, sponsor, PIPE investor) level. The dependent variable, MLOT, measures the money left on the table the PIPE investor makes from the SPAC deal. The key independent variables are the MLOT earned with the same sponsor in the last deal for all PIPE investors, premium PIPE investors, and non-premium PIPE investors. Additional controls include the log of IPO proceeds, the number of PIPE investors, the number of premium PIPE investors, days remaining, and separate fixed effects for the month, sponsor, and PIPE. All continuous variables are winsorized at the and percentile values. All other variables are defined in Table B1. Robust standard errors are reported in parentheses. , and denote p-values less than 0.1, 0.05, and 0.01, respectively.
@@
where is a binary variable equal to one if the PIPE (at) investor is a non-premium investor. The results from the estimation are in column (2) of Table 9. Comparing the second and third rows, we find that MLOT exhibits mean reversion for non-premium PIPE investors but not for premium investors. This result is consistent with the previous results that only the premium PIPE investors generate information, and Hypothesis 1-2.
@@
We use the instrumental variable (IV) approach to address this problem. The idea behind our choice of IV is based on the following observation: suppose is endogenous because is endogenous, then can be an IV for under appropriate additional assumptions. For a recent application of this intuition, see Fabra and Reguant (2014) and Aryal, Ciliberto and Leyden (2022). In our setting, is and is the Non-Premium dummy, and given the specification in (4) is uncorrelated with the quality of the current deal.
@@
The results from this estimation are in column (3) of Table 9. The first-stage Cragg-Donald Wald F-statistic of 92.82 suggests that our excluded variables are not weak instruments. Comparing these estimates with the ones in column (2), we see that, as expected, the fixed-effect estimate for the non-premium investor is biased upwards. More importantly, and consistent with the quid pro quo hypothesis, we find that one dollar higher mean deviation of MLOT in the last deal leads to a $2.729 reduction in the mean deviation from the current deal and vice versa presumably because now the non-premium investor is “asked” to participate in tepid deals that have smaller underpricing. In contrast, for the premium PIPE investors, the effect is the opposite: if the investor made one more dollar in MLOT above the mean in the last deal, on average, the MLOT in the next deal with the same sponsor increases by 23 cents consistent with the hypothesis that premium PIPE investors generate value-relevant information.131313Some deals have more than one sponsor. In such cases, we treated each sponsor separately. Although we do not show this here, we verify that the results are qualitatively the same if we exclude any SPAC with more than one sponsor or consider only the sponsor with the maximum number of past SPAC deals.

### v5

@@ 4 Data
We consider the universe of (1,072) SPACs publicly listed from January 1, 2010, to December 31, 2021. We follow these SPACs and record all major events and market returns until February 17, 2023. See Figure A2 for the monthly scatter plot in Figure A2, which includes the so-called boom and bust cycle for SPACs. We construct our sample by merging five datasets: (i) the SPAC Research,777https://www.spacresearch.com/. (ii) the PrivateRaise–SPACs Platform,888https://www.privateraise.com/about/platform_spac.php. (iii) the PrivateRaise–PIPEs Platform (iv) Bloomberg data on SPAC prices, and (v) Super 8-K filings from SEC.999https://www.sec.gov/edgar/search-and-access.,101010Thus, we exclude SPACs traded in the over-the-counter markets. After the merger between a SPAC and a target is complete, the SPAC is delisted. The new (merged) company may be assigned a new ticker symbol. We use Bloomberg-assigned unique identifiers to match across samples.
The SPAC Research and PrivateRaise’s SPACs platform databases contain information on SPACs, including important event dates. PrivateRaise’s PIPE platform database includes comprehensive information on PIPE transactions. The platform covers more than 20,000 placement profiles of PIPE transactions. It also provides contract terms, investment amounts, types of transactions, prices, identities, and allocations of the PIPE investors.
For these 1,072 SPACs, we manually merged the first two databases. The issuer’s name recorded in the PIPE transactions can be either that of SPAC or the post-merger company. We use both names to match the SPAC sample with the PrivateRaise database in those cases. We also use the total PIPE amount from PrivateRaise to cross-check the match to ensure match accuracy. If there are inconsistencies, we use the SPAC’s SEC filings documents to record the correct amount. We then merge our sample with Bloomberg’s pricing data; warrant prices are unavailable in CRSP. Our matching process yields detailed information on transactions for all three types of PIPE investments for each SPAC in our sample.
@@
Table 2 provides summary data. On average, SPACs raise $272.42 million proceeds during their IPO, with a median of $240 million and 90th percentile at $440 million. A typical SPAC unit contains 0.44 warrants and 0.01 rights. During the IPO process, SPACs’ sponsors exercise the over-allotment option (on average) 11.59
@@
Money Left on the Table. We construct a measure money left on the table (“MLOT”), which is the direct analog of the underpricing measure used in the traditional IPO literature. Specifically, we MLOT denotes the difference between the closing-day price on the first trading day111111As in the IPO literature, the motive for using first-day closing prices is weak-form efficiency, i.e., the first-day closing price reflects the best estimate of future cash flows using all available information. Longer holding period horizons would capture the same phenomena while introducing more noise into our tests. of the post-merger company and the offer price presented to PIPE investors by the SPAC, and this difference is multiplied by the number of securities held by the PIPE investors. This calculation considers both common shares and warrants acquired by the PIPE investors. To determine the rate of underpricing, we normalize the MLOT by PIPE investors’ total investment amount.121212Although the contracts with PIPE investors are signed before the merger is closed, they represent a pre-commitment price, implying that the PIPE investors deploy their capital only at the merger’s closure. This arrangement mirrors a ”one-day” return that traditional IPO investors experience. We carry out the MLOT and underpricing calculations for all three forms of PIPE investors (pre, at, and post).
@@
Premium PIPE Investors. In total, there are 2,005 PIPE investors in our sample. Out of them, 76 invested in PIPE (pre) transaction, 249 in PIPE (post), and 1,826 invested in at least one PIPE (at) transaction, but only a few large institutional investors participated frequently and invested significantly more than others. To capture this distinction, we use k-means clustering on the total investments in all SPACs and the number of SPACs to classify PIPE investors into two groups. We refer to a subset of PIPE investors who frequently participate as “premium” PIPE investors and the rest as non-premium PIPE investors. There are 34 premium PIPE investors in our sample (Table 3).
@@
This comparison is consistent with the quid pro quo arrangement where non-premium investors attain high returns on successful deals, offset by low returns on less successful ones. In contrast, premium investors see more steady returns due to their superior information.
@@ 5.1.1 Underpricing during SPAC IPO
We first examine underpricing during the SPAC IPO. Table 2 reports that the average equal-weighted first-day return of SPACs is 1.3%. Although our data start from 2010, most SPAC IPOs occurred between 2020 and 2021. In 2020, 248 SPACs went public, with an average equal-weighted first-day return of 1.3%; in 2021, 613 SPACs went public, and their average equal-weighted first-day return was 1.51%. Although these measures are statistically different from zero (p-values in both cases), they are small.131313The average equal-weighted first-day return for traditional IPOs was 41.6% in 2020 and 32% in 2021; see https://site.warrington.ufl.edu/ritter/files/IPOs-Underpricing.pdf (accessed 09/22/2022). In addition, SPACs almost always trade at $10 per share until the target announcement date. This uniformity is inconsistent with any information about sponsor quality revealed during the SPAC IPO.
@@ 5.1.2 Underpricing during de-SPAC
From Table 2, we see that, on average, a private company that uses the SPAC route to access the public market leaves $126.93 million MLOT to the sponsors and $201.22 million to the PIPE investors. Furthermore, we also see that PIPE investors who participate in different stages of SPAC are compensated differently. During the SPAC IPO, PIPE (pre) investors earn an average of $29.37 million through forward-purchase agreements, but PIPE (at) investors get, on average, a total of $193.07 million per deal. After the target is disclosed, PIPE (post) investors who participate receive an additional $8.07 million MLOT. Even among PIPE (at) investors, approximately one-third of the MLOT accrues to 34 premium investors.
Overall, the SPAC shares offered to PIPE (at) investors, especially the premium PIPE investors, during the de-SPAC phase are underpriced. These data features are consistent with PIPE (at) investors participating in the de-SPAC stage being privately informed about the quality of the target company.
5.2 Certification and Agency Theories
@@ 5.2.1 Redemption Rates
In this section, we examine Hypothesis 1-1 and Hypothesis 2-1 using the following regression:
where denotes the investment allocated to premium PIPE (at) investors as a fraction of the total IPO proceeds in SPAC , similarly for other allocations, and is a vector of covariates including total IPO proceeds and fixed effects that may vary across specifications. For example, if premium investors generate value-relevant information, the higher the share allocated to premium investors, the lower the redemption rate, i.e., . For inference, we cluster the robust standard errors at the sponsor level to capture correlations within a sponsor but across investors and SPACs.
Table 4 provides evidence consistent with both hypotheses. In column (1), we include only the allocations to four types of PIPE investors. In column (2), we control for the size of the SPAC IPO and the liquidation risk of the SPAC when the target is announced using Days Remaining until the expiration. In column (3), we control for the quarter-fixed effects. In column (4), we control for the target company’s sector fixed effects to alleviate the problem that investors may have varying investment preferences across industries.
@@
Note: This table shows the results from estimating (1) using SPACs that went through an IPO from January 2010 to December 2021. The independent variable is the investment amount as a percentage of the SPAC’s IPO proceeds for all PIPE (pre) investors, premium PIPE (at) investors, non-premium PIPE (at) investors, and all PIPE (post) investors. Premium PIPE (at) investors are listed in Table 3. All other variables are defined in Table B1. All unbounded continuous variables are winsorized at their and percentile values. Robust standard errors are clustered at the sponsor level and are reported in parentheses. In column (5), we use 1,000 Bootstrap samples to estimate the standard errors.. , and ∗∗∗ denote p-values less than 0.1, 0.05, and 0.01, respectively.
We find that the allocations to premium PIPE (at) investors negatively correlate with the redemption rates, and this result is robust across all specifications. In particular, the estimates suggest that one standard deviation (32.85 pps) increase in premium PIPE investors’ participation is associated with a 7.03 pps decrease in the redemption rate, which equals 12.4% of the sample mean. This evidence is consistent with both Hypothesis 1-1 and Hypothesis 2-1.
@@
The estimates with the additional controls are in Table 4 column (5). Comparing the estimates in columns (4) and (5), we can see that our result is qualitatively similar. The estimates suggest that one standard deviation (32.85 pps) increase in premium PIPE (at) investors’ participation is associated with a 7.19 pps decrease in the redemption rate, which equals 12.7% of the sample mean. This evidence is consistent with both Hypothesis 1-1 and Hypothesis 2-1.
The other coefficients in Table 4 are similarly intuitive. We also find no association between non-premium PIPE (at) or PIPE (pre) allocations and redemption rates. Thus, non-premium investors act neither in a certification role nor an agency role. Additionally, PIPE (pre) investors cannot certify an issue, as there is no target identified at that point. Finally, PIPE (post) loads positively in all specifications, reflecting instances in which PIPE investors are brought in to shore up funding shortfalls caused by redemption.
@@ 5.2.2 Announcement-day Return
Note: This table shows the results from estimating (1), announcement-day return as the dependent variable, using SPACs that went through an IPO from January 2010 to December 2021. The key independent variable is the investment amount as a percentage of the SPAC’s IPO proceeds for all PIPE (pre) investors, premium PIPE (at) investors, non-premium PIPE (at) investors, and all PIPE (post) investors. Premium PIPE (at) investors are listed in Table 3. All other variables are defined in Table B1. All unbounded continuous variables are winsorized at their and percentile values. Robust standard errors are clustered at the sponsor level and are reported in parentheses. In column (5), we use 1,000 Bootstrap samples to estimate the standard errors. , and ∗∗∗ denote p-values less than 0.1, 0.05, and 0.01, respectively.
@@
Likewise, the hypothesis predicts a negative market reaction to the announcement of PIPE (post) investments. We use an event study approach to assess this claim. In particular, we estimate the cumulative abnormal returns (CAR) (Campbell, Lo and MacKinlay, 1997) associated with announcements about PIPE investors while also announcing the target. The estimation period is from the SPAC’s unit separation date to the event window.
@@
Note: This table reports the SPAC share’s cumulative average abnormal return in percentage for three different event windows around announcements of 81 PIPE (post) transactions. Here, CAPM stands for capital asset pricing model, and FF3 and FF5 stand for Fama-French three- and five-factor models, respectively. The cross-sectional standard deviations of the cumulative abnormal returns are reported in parentheses. , and ∗∗∗ denote p-values less than 0.1, 0.05, and 0.01, respectively.
5.3 Insurance Theory
Next, we test Hypothesis 3 that there is a quid pro quo arrangement between sponsors and relationship investors. In particular, if the sponsor allocates securities to certain investors as a favor, and if the past MLOT is high, the sponsor will likely ask this investor to finance a weak deal today. We verify that the probability that a sponsor allocates securities to such a relationship investor increases if the investor has higher MLOT from the past deals and the SPAC deal is cold. Lastly, we show these relationships reduce the SPAC liquidation risks.
@@ 5.3.1 Mean-reverting MLOT
where MLOTp,s,i is the MLOT that PIPE (at) investor gets in the current SPAC deal initialized by the sponsor , and is the MLOT from the most recent deal in the past. As before, includes several covariates, including fixed effects (see Table 9).
@@
According to the certification hypothesis and agency theory, the deviation of lagged MLOT from its mean is either uncorrelated with the current MLOT; hence in (3). Under the insurance hypothesis, however, the MLOT should increase if the past deal was lower than the average and vice versa; hence, . The results are in column (1) of Table 9. Consistent with the insurance hypothesis, MLOT exhibits mean reversion.
@@
| (lag-MLOT-)(lag2-MLOT-) | 121.96 | 85.54 | |
| (lag2-MLOT-)2 | 135.78 | 5.77 | |
| (lag-MLOT-)2 | 166.93 | 297.54 | |
Note: This table shows results from estimating (4) using SPACs that went through IPO between January 2010 to December 2021. Each observation is at (SPAC, sponsor, PIPEat investor) level. The dependent variable, MLOT, measures the money left on the table the PIPEat investor makes from the SPAC deal. The key independent variables are the MLOT earned with the same sponsor in the last deal for all PIPEat investors, premium PIPEat investors, and non-premium PIPEat investors. Additional controls include the log of IPO proceeds, the number of PIPE investors, the number of premium PIPE investors, days remaining, and separate fixed effects for the month, sponsor, and PIPE. All continuous variables are winsorized at the and percentile values. All other variables are defined in Table B1. Robust standard errors are reported in parentheses. , and ∗∗∗ denote p-values less than 0.1, 0.05, and 0.01, respectively.
@@
where is a binary variable equal to one if the PIPE (at) investor is a non-premium investor. The results from the estimation are in column (2) of Table 9. Comparing the second and third rows, we find that MLOT exhibits mean reversion for non-premium PIPE investors but not for premium investors. Synthesizing with our prior results, we conclude that premium investors act in a certification role or earn agency rents, whereas non-premium investors serve in an insurance role.
@@
We use the instrumental variable (IV) approach to address this problem. The idea behind our choice of IV is based on the following observation: suppose is endogenous because is endogenous, then can be an IV for under additional assumptions. For a recent application of this intuition, see Fabra and Reguant (2014) and Aryal, Ciliberto and Leyden (2022). In our setting, is and is the Non-Premium dummy, and is uncorrelated with the quality of the deal.
@@
The results from this estimation are in column (3) of Table 9. The first-stage Cragg-Donald Wald F-statistic of 92.82 suggests that our excluded variables are not weak instruments. Comparing these estimates with the ones in column (2), we see that, as expected, the fixed-effect estimate for the non-premium investor is biased upwards. More importantly, and consistent with the quid pro quo hypothesis, we find that one dollar higher mean deviation of MLOT in the last deal leads to a $2.729 reduction in the mean deviation from the current deal and vice versa presumably because now the non-premium investor is “asked” to participate in tepid deals that have smaller underpricing. In contrast, for the premium PIPE investors, the effect is the opposite: if the investor made one more dollar in MLOT above the mean in the last deal, on average, the MLOT in the next deal with the same sponsor increases by 23 cents consistent with the hypothesis that premium PIPE investors generate value-relevant information.141414If a deal has more than one sponsor, we treat each sponsor separately. We verify that the results are qualitatively similar if we exclude SPACs with multiple sponsors.